# Turning Waste into Wealth: Leveraging Low-Quality Samples for Enhancing Continuous Conditional Generative Adversarial Networks

### 14 Dec 2023 v2

@@ Introduction
The objective of Continuous Conditional Generative Modeling (CCGM), as illustrated in Fig. 2, is to estimate the distribution of high-dimensional data, such as images, in terms of continuous scalar variables, referred to as regression labels. However, this task is very challenging due to insufficient or even zero training images for certain regression labels and the absence of a suitable label input mechanism.
@@
To alleviate the data sparsity or imbalance issue, traditional data augmentation techniques for GANs (Zhao et al. 2020; Karras et al. 2020; Tran et al. 2021; Jiang et al. 2021; Tseng et al. 2021; Liu et al. 2023) often employ geometric transformations on real images, such as flipping, translation, and rotation, to guide GANs in learning “what to generate”. However, Sinha et al. (2021) introduced a distinctive approach, known as Negative Data Augmentation (NDA), for unconditional or class-conditional GANs. The approach, depicted in Fig. 3, intentionally crafts negative samples via transformations applied to training images, encompassing techniques like Jigsaw (Noroozi and Favaro 2016), Stitching, Mixup (Zhang et al. 2018), Cutout (DeVries and Taylor 2017), and CutMix (Yun et al. 2019). These negative samples, akin to those produced by the generator, are presented as fake images (in contrast to real images from the training set) and incorporated into the discriminator’s training, instructing GANs on “what to avoid”. Nevertheless, NDA’s application is limited in the context of CcGANs, as it cannot replicate negative samples that may arise from pre-trained CcGANs, illustrated by two types of representative low-quality images in Fig. 4.
To address this challenge, we introduce a novel NDA strategy termed Dual-NDA in this work, as depicted in Fig. 1. Unlike the synthetic images showcased in Fig. 3, Dual-NDA enriches the training set of CcGANs with two categories of negative samples, strategically mirroring the low-quality images seen in Fig. 4. Type I negative samples are mismatched image-label pairs formed by manipulating labels of real images in the training set. In contrast, Type II negative samples are generated by evaluating fake images from a pre-trained generator and retaining those exhibiting the poorest visual quality. Based on these two types of negative samples, we devise a new soft vicinal discriminator loss tailored to enhance the training of CcGANs. Our comprehensive experimental investigation demonstrates the effectiveness of these negative samples in improving CcGAN performance, particularly in terms of visual quality and label consistency enhancement. Notably, CcGANs can generally achieve remarkable superiority through Dual-NDA over state-of-the-art class-conditional GANs and diffusion models.
Fig. 1: Illustrative workflows for the vanilla NDA (Sinha et al. 2021) and our proposed Dual-NDA.
Fig. 2: Illustration of the CCGM task and sample images from two regression datasets (UTKFace and Steering Angle).
Fig. 3: Example negative samples from NDA by transforming a realistic training image (from Sinha et al. (2021)).
Fig. 4: Some actual low-quality fake images generated from a pre-trained CcGAN at “Age=3” on UTKFace. The term “label-inconsistent” indicates that these fake images do not align with the conditioning label.
@@
To enhance unconditional and class-conditional GANs, Sinha et al. (2021) introduced a novel Negative Data Augmentation (NDA) strategy that intentionally creates out-of-distribution negative samples by applying various transformations, such as Jigsaw, Stitching, and Mixup, to realistic training images. Furthermore, Sinha et al. (2021) introduced a new GAN training objective incorporating these negative samples. This training objective guides the generator network away from producing low-quality outputs resembling the negative samples, thereby encouraging the generator to create more realistic and diverse images. An illustrative workflow of NDA is provided in Fig. 1. NDA has demonstrated its effectiveness in enhancing the performance of BigGAN (Brock, Donahue, and Simonyan 2019), a popular cGAN model, across both unconditional and class-conditional generative modeling tasks, highlighting NDA’s potential for advancing GAN-driven image synthesis. However, as demonstrated in Fig. 3 and Fig. 4, NDA-generated negative samples markedly differ from the low-quality samples produced by CcGANs, which explains the limited impact of NDA in our experimental study. Consequently, we propose a novel NDA strategy named Dual-NDA, aimed at generating the two distinct types of negative samples illustrated in Fig. 4.
@@
We present Dual-NDA, an innovative NDA strategy tailored to enhance the performance of CcGANs. The comprehensive workflow of Dual-NDA is depicted in Fig. 1. This approach introduces a novel training paradigm for CcGANs, comprising two key components: a carefully designed negative sample generation mechanism adept at imitating the low-quality images shown in Fig. 4, and a new vicinal discriminator loss that harnesses these negative samples to enhance the visual quality and label consistency of generated images.
@@ A Modified Training Mechanism for CcGANs
Datasets. Following Ding et al. (2023b), we utilize two regression datasets, namely UTKFace (Zhang, Song, and Qi 2017) and Steering Angle (Chen 2018), for our experimental study. The UTKFace dataset includes human facial images annotated with ages for regression analysis. We use the pre-processed version (Ding et al. 2023b), which consists of 14,723 RGB images. Age labels span from 1 to 60, showcasing a broad age distribution with varying sample sizes per age group (ranging from 50 to 1051 images). The Steering Angle dataset, derived from a subset of an autonomous driving dataset (Chen 2018), comprises 12,271 RGB images. These images are annotated with 1,774 distinct steering angles, spanning from $-80.0^{\circ}$ to $80.0^{\circ}$. All Steering Angle images are captured using a dashboard camera on a car, with simultaneous recording of the corresponding steering wheel rotation angles. Both datasets offer two versions, encompassing resolutions of $64\times 64$ and $128\times 128$, respectively, resulting in a total of four datasets. Illustrative training images are visualized in Fig. 2, while data distributions are presented in Appendix B.
@@
We present a performance comparison of various methods on the four datasets in Table 1, complemented by illustrative figures presented in Fig. 5 and Fig. 6. Analysis of the table and figures reveal the following findings:
@@
  • Compared with the baseline “CcGAN w/o NDA”, Dual-NDA exhibits reduced values in terms of NIQE and Label Score metrics. This trend is further supported by Fig. 5 and Fig. 6, which illustrate a substantial reduction in NIQE and Label Score values across the entire range of regression labels. These outcomes suggest that incorporating Type I and II negative samples effectively enhances the visual quality and label consistency of the generated samples.
@@
[…]
Fig. 5: Line graphs of NIQE/Label Score versus Age for the UTKFace (128$\times$128) experiment.
Fig. 6: Line graphs of NIQE/Label Score versus Age for the Steering Angle (128$\times$128) experiment.
@@
We also conduct comprehensive ablation studies on the Steering Angle (64$\times$64) dataset. These studies are designed to systematically investigate the impact of individual components and key hyperparameters within the framework of Dual-NDA, as outlined below.
@@
    The Effects of Type I and II Negative Samples: We individually delve into the distinct impacts of Type I and Type II negative samples. The experimental results are summarized in Fig. 7. From this visual result, we can deduce that the integration of Type I negative samples during training effectively enhances the label consistency of generated images (evidenced by decreased Label Score values), yet it might potentially lead to a compromise in visual quality (reflected by higher NIQE values relative to the blue dashed line). Interestingly, the influence of Type II negative samples exhibits a unique pattern. These samples, designed to enhance visual quality, surprisingly contribute to an improvement in label consistency (as reflected by a Label Score value below the red dashed line). This phenomenon can be rationalized by the insights provided by Ding et al. (2023a), which suggest that certain label-inconsistent fake images often exhibit poor visual quality. Consequently, the removal of visually unrealistic images may lead to an enhancement in label consistency. Combining both Type I and Type II negative samples ultimately leads to the most favorable overall performance.
@@
    The Combined Impact of $\lambda_{1}$ and $\lambda_{2}$: Instead of examining the individual effects of $\lambda_{1}$ and $\lambda_{2}$, we delve into their combined impact on the SFID scores of CcGANs. We achieve this by varying $\lambda_{1}+\lambda_{2}$ within the range of 0.2 to 0.7. The experimental outcomes, shown in Fig. 8, highlight a consistent trend: Dual-NDA consistently outperforms the baseline CcGAN across a wide spectrum of $\lambda_{1}+\lambda_{2}$ values in terms of SFID. Notably, adjustments to $\lambda_{1}+\lambda_{2}$ within the interval of $[0.2,0.5]$ often yield optimal performance for CcGANs.
@@
    The Effect of $q_{2}$: In the last ablation study, we examine the influence of $q_{2}$ on Dual-NDA’s performance and report the quantitative results in Tab. 3. These results reveal that Dual-NDA is insensitive to the values of $q_{2}$, so we let $q_{2}=0.9$ throughout our experiments.
@@
Fig. 7: The individual impacts of Type I and Type II negative samples on Dual-NDA in the Steering Angle (64$\times$64) experiment. The SFID scores for compared settings are shown in parentheses. The blue and red dashed lines represent the NIQE and Label Score of “CcGAN w/o NDA”, respectively.
Fig. 8: The effect of $\lambda_{1}+\lambda_{2}$ on Dual-NDA in the Steering Angle (64$\times$64) experiment. The gray dashed line represents the SFID score of “CcGAN w/o NDA”.
@@
$\bm{q_{1}}$ SFID NIQE Diversity Label Score w/o NDA 1.334 (0.531) 1.784 (0.065) 1.234 (0.209) 14.807 (14.297) 0.3 1.125 (0.510) 1.759 (0.047) 1.271 (0.181) 11.365 (11.583) 0.5 1.114 (0.503) 1.738 (0.055) 1.251 (0.172) 11.809 (11.694) 0.7 1.161 (0.540) 1.763 (0.043) 1.227 (0.179) 11.162 (11.241) 0.9 1.187 (0.558) 1.763 (0.045) 1.270 (0.169) 11.550 (11.163)
@@
$\bm{q_{2}}$ SFID NIQE Diversity Label Score w/o NDA 1.334 (0.531) 1.784 (0.065) 1.234 (0.209) 14.807 (14.297) 0.5 1.107 (0.506) 1.717 (0.058) 1.301 (0.170) 11.881 (11.144) 0.6 1.193 (0.592) 1.733 (0.045) 1.278 (0.205) 11.459 (11.675) 0.7 1.150 (0.494) 1.763 (0.038) 1.278 (0.158) 13.161 (13.346) 0.8 1.174 (0.563) 1.756 (0.040) 1.254 (0.186) 11.207 (11.231) 0.9 1.114 (0.503) 1.738 (0.055) 1.251 (0.172) 11.809 (11.694)
Table 4: Generating Type II negative samples via using different generators.
Generator SFID $\downarrow$ NIQE $\downarrow$ Diversity $\uparrow$ Label Score $\downarrow$ None 1.334 (0.531) 1.784 (0.065) 1.234 (0.209) 14.807 (14.297) ADCGAN 1.212 (0.481) 1.832 (0.055) 1.222 (0.180) 10.884 (11.041) ReACGAN 1.203 (0.499) 1.800 (0.078) 1.273 (0.171) 11.486 (10.904) CcGAN 1.114 (0.503) 1.738 (0.055) 1.251 (0.172) 11.809 (11.694)
@@ B.2 The Data Distribution of UTKFace and Steering Angle
All samples from both the UTKFace and Steering Angle datasets are employed for training purposes. The distribution of these samples is depicted in Fig. 9, revealing pronounced data imbalance issues within both datasets. This issue becomes more severe in the Steering Angle datasets, where some steering angle values are only represented by less than 10 images.
Fig. 9: The data distribution of the UTKFace and Steering Angle datasets.
@@ B.3 Detailed Training Setups
Table 5: Important training setups for baseline methods in Table 1. Please note that, for implementing Dual-NDA, we utilize the pre-trained generator of “CcGAN w/o NDA” in conjunction with the NIQE filtering process to generate Type II negative samples.
@@ Appendix C Extra Experimental Results
C.1 Some extra ablation studies

### 31 Dec 2023 v3

@@ Introduction
The objective of Continuous Conditional Generative Modeling (CCGM), as illustrated in Figure 2, is to estimate the distribution of high-dimensional data, such as images, in terms of continuous scalar variables, referred to as regression labels. However, this task is very challenging due to insufficient or even zero training images for certain regression labels and the absence of a suitable label input mechanism.
@@
To alleviate the data sparsity or imbalance issue, traditional data augmentation techniques for GANs (Zhao et al. 2020; Karras et al. 2020; Tran et al. 2021; Jiang et al. 2021; Tseng et al. 2021; Liu et al. 2023) often employ geometric transformations on real images, such as flipping, translation, and rotation, to guide GANs in learning “what to generate”. However, Sinha et al. (2021) introduced a distinctive approach, known as Negative Data Augmentation (NDA), for unconditional or class-conditional GANs. The approach, depicted in Figure 3, intentionally crafts negative samples via transformations applied to training images, encompassing techniques like Jigsaw (Noroozi and Favaro 2016), Stitching, Mixup (Zhang et al. 2018), Cutout (DeVries and Taylor 2017), and CutMix (Yun et al. 2019). These negative samples, akin to those produced by the generator, are presented as fake images (in contrast to real images from the training set) and incorporated into the discriminator’s training, instructing GANs on “what to avoid”. Nevertheless, NDA’s application is limited in the context of CcGANs, as it cannot replicate negative samples that may arise from pre-trained CcGANs, illustrated by two types of representative low-quality images in Figure 4.
To address this challenge, we introduce a novel NDA strategy termed Dual-NDA in this work, as depicted in Figure 1. Unlike the synthetic images showcased in Figure 3, Dual-NDA enriches the training set of CcGANs with two categories of negative samples, strategically mirroring the low-quality images seen in Figure 4. Type I negative samples are mismatched image-label pairs formed by manipulating labels of real images in the training set. In contrast, Type II negative samples are generated by evaluating fake images from a pre-trained generator and retaining those exhibiting the poorest visual quality. Based on these two types of negative samples, we devise a new soft vicinal discriminator loss tailored to enhance the training of CcGANs. Our comprehensive experimental investigation demonstrates the effectiveness of these negative samples in improving CcGAN performance, particularly in terms of visual quality and label consistency enhancement. Notably, CcGANs can generally achieve remarkable superiority through Dual-NDA over state-of-the-art class-conditional GANs and diffusion models.
Figure 1: Illustrative workflows for the vanilla NDA (Sinha et al. 2021) and our proposed Dual-NDA.
Figure 2: Illustration of the CCGM task and sample images from two regression datasets (UTKFace and Steering Angle).
Figure 3: Example negative samples from NDA by transforming a realistic training image (from Sinha et al. (2021)).
Figure 4: Some actual low-quality fake images generated from a pre-trained CcGAN at “Age=3” on UTKFace. The term “label-inconsistent” indicates that these fake images do not align with the conditioning label.
@@
To enhance unconditional and class-conditional GANs, Sinha et al. (2021) introduced a novel Negative Data Augmentation (NDA) strategy that intentionally creates out-of-distribution negative samples by applying various transformations, such as Jigsaw, Stitching, and Mixup, to realistic training images. Furthermore, Sinha et al. (2021) introduced a new GAN training objective incorporating these negative samples. This training objective guides the generator network away from producing low-quality outputs resembling the negative samples, thereby encouraging the generator to create more realistic and diverse images. An illustrative workflow of NDA is provided in Figure 1. NDA has demonstrated its effectiveness in enhancing the performance of BigGAN (Brock, Donahue, and Simonyan 2019), a popular cGAN model, across both unconditional and class-conditional generative modeling tasks, highlighting NDA’s potential for advancing GAN-driven image synthesis. However, as demonstrated in Figure 3 and Figure 4, NDA-generated negative samples markedly differ from the low-quality samples produced by CcGANs, which explains the limited impact of NDA in our experimental study. Consequently, we propose a novel NDA strategy named Dual-NDA, aimed at generating the two distinct types of negative samples illustrated in Figure 4.
@@
We present Dual-NDA, an innovative NDA strategy tailored to enhance the performance of CcGANs. The comprehensive workflow of Dual-NDA is depicted in Figure 1. This approach introduces a novel training paradigm for CcGANs, comprising two key components: a carefully designed negative sample generation mechanism adept at imitating the low-quality images shown in Figure 4, and a new vicinal discriminator loss that harnesses these negative samples to enhance the visual quality and label consistency of generated images.
@@ A Modified Training Mechanism for CcGANs
Datasets. Following Ding et al. (2023b), we utilize two regression datasets, namely UTKFace (Zhang, Song, and Qi 2017) and Steering Angle (Chen 2018), for our experimental study. The UTKFace dataset includes human facial images annotated with ages for regression analysis. We use the pre-processed version (Ding et al. 2023b), which consists of 14,723 RGB images. Age labels span from 1 to 60, showcasing a broad age distribution with varying sample sizes per age group (ranging from 50 to 1051 images). The Steering Angle dataset, derived from a subset of an autonomous driving dataset (Chen 2018), comprises 12,271 RGB images. These images are annotated with 1,774 distinct steering angles, spanning from $-80.0^{\circ}$ to $80.0^{\circ}$. All Steering Angle images are captured using a dashboard camera on a car, with simultaneous recording of the corresponding steering wheel rotation angles. Both datasets offer two versions, encompassing resolutions of $64\times 64$ and $128\times 128$, respectively, resulting in a total of four datasets. Illustrative training images are visualized in Figure 2, while data distributions are presented in Appendix B.
@@
We present a performance comparison of various methods on the four datasets in Table 1, complemented by illustrative figures presented in Figure 5 and Figure 6. Analysis of the table and figures reveal the following findings:
@@
  • Compared with the baseline “CcGAN w/o NDA”, Dual-NDA exhibits reduced values in terms of NIQE and Label Score metrics. This trend is further supported by Figure 5 and Figure 6, which illustrate a substantial reduction in NIQE and Label Score values across the entire range of regression labels. These outcomes suggest that incorporating Type I and II negative samples effectively enhances the visual quality and label consistency of the generated samples.
@@
Figure 5: Line graphs of NIQE/Label Score versus Age for the UTKFace (128$\times$128) experiment.
Figure 6: Line graphs of NIQE/Label Score versus Angle for the Steering Angle (128$\times$128) experiment.
@@
We also conduct comprehensive ablation studies on the Steering Angle (64$\times$64) dataset. These studies are designed to systematically investigate the impact of individual components and key hyperparameters within the framework of Dual-NDA, as outlined below. The standard deviations for the evaluation scores are provided in Appendix C.
@@
    The Effects of Type I and II Negative Samples: We individually delve into the distinct impacts of Type I and Type II negative samples. The experimental results are summarized in Figure 7. From this visual result, we can deduce that the integration of Type I negative samples during training effectively enhances the label consistency of generated images (evidenced by decreased Label Score values), yet it might potentially lead to a compromise in visual quality (reflected by higher NIQE values relative to the blue dashed line). Interestingly, the influence of Type II negative samples exhibits a unique pattern. These samples, designed to enhance visual quality, surprisingly contribute to an improvement in label consistency (as reflected by a Label Score value below the red dashed line). This phenomenon can be rationalized by the insights provided by Ding et al. (2023a), which suggest that certain label-inconsistent fake images often exhibit poor visual quality. Consequently, the removal of visually unrealistic images may lead to an enhancement in label consistency. Combining both Type I and Type II negative samples ultimately leads to the most favorable overall performance.
@@
    The Combined Impact of $\lambda_{1}$ and $\lambda_{2}$: Instead of examining the individual effects of $\lambda_{1}$ and $\lambda_{2}$, we delve into their combined impact on the SFID scores of CcGANs. We achieve this by varying $\lambda_{1}+\lambda_{2}$ within the range of 0.2 to 0.7. The experimental outcomes, shown in Figure 8, highlight a consistent trend: Dual-NDA consistently outperforms the baseline CcGAN across a wide spectrum of $\lambda_{1}+\lambda_{2}$ values in terms of SFID. Notably, adjustments to $\lambda_{1}+\lambda_{2}$ within the interval of $[0.2,0.5]$ often yield optimal performance for CcGANs.
@@
    The Effect of $q_{2}$: In the last ablation study, we examine the influence of $q_{2}$ on Dual-NDA’s performance and report the quantitative results in Table 3. These results reveal that Dual-NDA is insensitive to the values of $q_{2}$, so we let $q_{2}=0.9$ throughout our experiments.
@@
Figure 7: The individual impacts of Type I and Type II negative samples on Dual-NDA in the Steering Angle (64$\times$64) experiment. The SFID scores for compared settings are shown in parentheses. The blue and red dashed lines represent the NIQE and Label Score of “CcGAN w/o NDA”, respectively.
Figure 8: The effect of $\lambda_{1}+\lambda_{2}$ on Dual-NDA in the Steering Angle (64$\times$64) experiment. The gray dashed line represents the SFID score of “CcGAN w/o NDA”.
$\bm{q_{1}}$ SFID NIQE Diversity Label Score
w/o NDA 1.334 1.784 1.234 14.807
0.3 1.125 1.759 1.271 11.365
0.5 1.114 1.738 1.251 11.809
0.7 1.161 1.763 1.227 11.162
0.9 1.187 1.763 1.270 11.550
@@
$\bm{q_{2}}$ SFID NIQE Diversity Label Score
w/o NDA 1.334 1.784 1.234 14.807
0.5 1.107 1.717 1.301 11.881
0.6 1.193 1.733 1.278 11.459
0.7 1.150 1.763 1.278 13.161
0.8 1.174 1.756 1.254 11.207
0.9 1.114 1.738 1.251 11.809
@@
Generator SFID NIQE Diversity Label Score
None 1.334 1.784 1.234 14.807
ADCGAN 1.212 1.832 1.222 10.884
ReACGAN 1.203 1.800 1.273 11.486
CcGAN 1.114 1.738 1.251 11.809
Table 4: Type II negative samples from different generators.
@@ B.2 The Data Distribution of UTKFace and Steering Angle
All samples from both the UTKFace and Steering Angle datasets are employed for training purposes. The distribution of these samples is depicted in Figure 9, revealing pronounced data imbalance issues within both datasets. This issue becomes more severe in the Steering Angle datasets, where some steering angle values are only represented by less than 10 images.
Figure 9: The data distribution of the UTKFace and Steering Angle datasets.
@@ B.3 Detailed Training Setups
Table 5: Important training setups for baseline methods in Table 1. Please note that, for Dual-NDA, we utilize the pre-trained generator of “CcGAN w/o NDA” in conjunction with the NIQE filtering process to generate Type II negative samples.
@@ Appendix C Extra Experimental Results
C.1 The complete results for Tables 2, 3, and 4 in the paper.
In the paper, owing to space constraints, we do not display the standard deviations for the evaluation scores in Tables 2, 3, and 4. Here, we present the comprehensive results, which are fully detailed in Table 6, Table 7, and Table 8, respectively.
$\bm{q_{1}}$ SFID NIQE Diversity Label Score
w/o NDA 1.334 (0.531) 1.784 (0.065) 1.234 (0.209) 14.807 (14.297)
0.3 1.125 (0.510) 1.759 (0.047) 1.271 (0.181) 11.365 (11.583)
0.5 1.114 (0.503) 1.738 (0.055) 1.251 (0.172) 11.809 (11.694)
0.7 1.161 (0.540) 1.763 (0.043) 1.227 (0.179) 11.162 (11.241)
0.9 1.187 (0.558) 1.763 (0.045) 1.270 (0.169) 11.550 (11.163)
$\bm{q_{2}}$ SFID NIQE Diversity Label Score
w/o NDA 1.334 (0.531) 1.784 (0.065) 1.234 (0.209) 14.807 (14.297)
0.5 1.107 (0.506) 1.717 (0.058) 1.301 (0.170) 11.881 (11.144)
0.6 1.193 (0.592) 1.733 (0.045) 1.278 (0.205) 11.459 (11.675)
0.7 1.150 (0.494) 1.763 (0.038) 1.278 (0.158) 13.161 (13.346)
0.8 1.174 (0.563) 1.756 (0.040) 1.254 (0.186) 11.207 (11.231)
0.9 1.114 (0.503) 1.738 (0.055) 1.251 (0.172) 11.809 (11.694)
Generator SFID $\downarrow$ NIQE $\downarrow$ Diversity $\uparrow$ Label Score $\downarrow$
None 1.334 (0.531) 1.784 (0.065) 1.234 (0.209) 14.807 (14.297)
ADCGAN 1.212 (0.481) 1.832 (0.055) 1.222 (0.180) 10.884 (11.041)
ReACGAN 1.203 (0.499) 1.800 (0.078) 1.273 (0.171) 11.486 (10.904)
CcGAN 1.114 (0.503) 1.738 (0.055) 1.251 (0.172) 11.809 (11.694)
Table 8: Type II negative samples from different generators.
C.2 Some extra ablation studies